Directrix (directrix):

If it takes 2 programmers 2 days to write 2 programs, how many programmers are needed to write 20 programs in 20 days?

OpenStudy (anonymous):

2 programs are written in 2 days by = 2 programmers 2 programs will be written in 1 day by = 2*2 = 4 programmers 1 program will be written in 1 day by = 4/2 = 2 programmers 20 programs will be written in 1 day by = 2*20 = 40 programmers 20 programs will be written in 20 days by = 40/20 = 2 programmers answer 2 programmers.

Directrix (directrix):

Anybody else have a solution. If so, post.

OpenStudy (anonymous):

@geniusy_2006 , looks good... i agree.

OpenStudy (ash2326):

Yeah it's correct, good illustration @geniusy_2006

OpenStudy (hoblos):

it is a great explanation @geniusy_2006 but you can think in an easier way though simply 2 programmers write 2 programs in 2 days so in 20 days they write [ 20*2/2 ] = 20 programs you get the answer is two